# Curvature fluctuations in baby quantum gravity models

### v1

@@ III Renormalization
Mirroring the procedure in [10], we may renormalize away $\mathcal{L}$. As both $\langle R_{av}\rangle$ and $\langle R_{av}^{2}\rangle$ are of physical significance, we focus on those. As, to leading order, $\langle k_{0}\rangle$ is proportional to $\mathcal{L}$, we use $\langle k_{0}\rangle$ to renormalize.
Note that physical quantities are calculated for a fixed $\mathcal{L}^{phys}$, but we will renormalize by dividing through by $G$, so that $\mathcal{L}=\mathcal{L}^{phys}/G$ and $\varepsilon=\varepsilon^{phys}/G$. Observe that all of the expressions in the previous sections are written in terms of physical quantities. Further, we can define $\langle k_{0}\rangle_{\mathcal{L}}=\langle k_{0}^{phys}\rangle/G$, $\langle R_{av}\rangle_{\mathcal{L}}=G\langle R_{av}^{phys}\rangle$, and $\langle R_{av}^{2}\rangle_{\mathcal{L}}=G^{2}\langle(R_{av}^{2})^{phys}\rangle$.
In particular, we have
$\langle k_{0}\rangle_{\mathcal{L}}\approx$ $\frac{3\mathcal{L}}{5}+\frac{3i}{160}$
$\langle R_{av}\rangle_{\mathcal{L}}\approx$ $\frac{9i}{8\mathcal{L}}-\frac{9i(144\sqrt{\pi}+35)}{2240\sqrt{\pi}\mathcal{L}^{5/2}}$
$\langle R_{av}^{2}\rangle_{\mathcal{L}}\approx$ $\frac{9}{64\sqrt{\pi}}\sqrt{\frac{1}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81}{1024\sqrt{\pi}}\sqrt{\frac{1}{\varepsilon^{2}\mathcal{L}^{7}}}\,.$
To perform renormalization, observe that
$\langle R^{phys}_{av}\rangle=\langle R_{av}\rangle_{\mathcal{L}}\frac{\langle k_{0}\rangle_{\mathcal{L}}}{\langle k_{0}^{phys}\rangle}\,.$
Now using the above expressions and then solving for $\mathcal{L}$ in terms of $\langle k_{0}^{phys}\rangle$, we find that
$\langle R^{phys}_{av}\rangle=\frac{81iG}{\langle k_{0}^{phys}\rangle^{2}}+\mathcal{O}(G^{3/2})\,.$
As expected, any non-vanishing behavior in the averaged Ricci scalar is purely quantum in nature.
Using the same method, we find that the renormalized variance of the Ricci scalar takes the form
$\langle(R^{2}_{av})^{phys}\rangle=\frac{81\sqrt{15}}{8000\sqrt{\pi}}\sqrt{\frac{G^{3}}{(\varepsilon^{phys})^{2}\langle k_{0}^{phys}\rangle^{5}}}+\mathcal{O}(G^{5/2})\,.$
Under renormalization, we see that this is a genuine quantum effect and that the behavior (if not the coefficient) remains stable under renormalization. Namely, a UV divergence still arises. Thus we may treat the behavior in Section II as corresponding to physical observables as well, and so it is appropriate to provide a physical interpretation of these results.
@@ IV Physical Interpretation
The construction in Section II provides a more detailed account of the behavior leading to large fluctuations than in [2, 3]. There, a fixed-size region of spacetime is specified and fluctuations are viewed as resonances in that volume, with the magnitude of the fluctuations depending on the scale of the spacetime volume. However, while this heuristic description is a useful starting point, the notion that one begins by considering a fixed volume spacetime is contrary to the spirit of a background-independent theory of quantum gravity. The model provided here using quantum geometry gives a properly background-independent approach that leads to curvature fluctuations.
Physically, we view this model as approximating the behavior of a “small” subset of the spacetime manifold prior to specification of the metric. By then imposing a quantum geometry on the manifold and minimum and maximum cutoffs on the sizes of the dimensions of that subset, we find that the approximate size of the subset is determined by the expectation value $\langle k_{0}\rangle$, inducing a metric. Unlike the heuristic in [2, 3], the magnitude of curvature fluctuations depends also on the minimum size of the region of spacetime. One way to see this is that derivatives of metric functions depend both on the discrete difference in the metric functions and inversely on the metric functions themselves. If the metric functions are allowed to approach zero, fluctuations can diverge.
Nonetheless, the behavior of this model still reproduces the hierarchy of energy densities required by [5] so long as appropriately physical regulators are introduced. As we intend to model only a “unit” of quantum spacetime in a way that is insensitive to matter field fluctuations, the IR cutoff should be shorter than the scale where standard model fields can see the fluctuations.
@@
These estimations support the critical assertion in [5] that $\rho_{QG}\gg\rho_{SM}$, implying that for a given region of spacetime the size of the Lorentzian square examined here, the scale factor is oscillatory rather than exponential. Furthermore, taking the IR length scale at $2\times 10^{-21}\text{ m}$, we find from $\langle k_{0}\rangle$ that the average size of the spacetime domain is on the order of $10^{-21}\text{ m}$, meaning that the wavelength of the fluctuations are of the same order. Thus, averaging over even a cubic centimeter of space would mean summing the fluctuations of $10^{56}$ microscopic domains of spacetime, clearly sufficiently many to average out fluctuations. Similarly, the time scale of these fluctuations is on the order of $10^{-30}\text{ s}$, whereas the smallest time yet measured [13] is on the order of $10^{-19}\text{ s}$, meaning that we may average over $10^{10}$ spacetime domains temporally—clearly sufficient to again average out the fluctuations. The implication is that, absent any other effects aside from the quantum geometric effects and the standard model vacuum fluctuations, the scale factor would be measured to be constant over any macroscopic scale.
While the above computation should be treated as qualitative, it does show that the argument used in [4, 5]—that the gravitational fluctuations dominate the standard model energy density when we consider the quantum gravity regime—is robust across many quantum gravitational models, even those as simple as a $1+1$d discrete quantum system as discussed above.
@@ V A Perturbative Cosmological Constant
In [5], it was also assumed that the intrinsic macroscopic cosmological constant of the spacetime also vanishes. Interestingly, the model constructed in Section II also provides a hint that this constraint is reasonable. To see this, we consider the behavior of curvature correlators in the presence of a small cosmological constant (such that we may keep only terms linear in $\Lambda$). To that end, we perturb the action and the corresponding correlators merely by writing
$Z_{\Lambda}=4\int_{-1}^{1}\int_{-1}^{1}\int_{\varepsilon}^{\mathcal{L}}\int_{\varepsilon}^{\mathcal{L}}{\rm d}k\,{\rm d}\ell\,{\rm d}k_{0}\,{\rm d}\ell_{0}\,k_{0}\ell_{0}\operatorname{exp}\left[\frac{i}{G}(S_{g}-4\Lambda k_{0}\ell_{0})\right]$
@@
$Z_{\Lambda}\approx\left(1-\tfrac{4i\Lambda}{G}|\langle k_{0}\rangle_{0}|^{2}\right)Z_{0}\,.$
@@
$=\left(1+\tfrac{4i\Lambda}{G}|\langle k_{0}\rangle_{0}|^{2}\right)\langle\mathcal{O}{\rangle}_{0}-\tfrac{4i\Lambda}{G}\langle k_{0}\ell_{0}\mathcal{O}\rangle_{0}\,,$
@@
$\langle k_{0}\ell_{0}R_{av}\rangle$ $\approx\frac{iG}{8}+\frac{27i}{2240}\sqrt{\frac{G^{5}}{\mathcal{L}^{3}}}$
$\langle k_{0}\ell_{0}R_{av}^{2}\rangle$ $\approx(0.0079+0.0264\ln G/\epsilon)(G/\mathcal{L})^{3/2}$
@@
$\langle R_{av}\rangle_{\Lambda}\approx-\frac{28}{25}\Lambda+0.95\Lambda\sqrt{\frac{G}{\mathcal{L}}}\,,$
@@
$\langle R_{av}^{2}\rangle_{\Lambda}\approx\frac{9}{64\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81i\Lambda}{400\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}}}.$
Note that, as expected, the introduction of a cosmological constant induces a non-vanishing Ricci scalar. However, $\langle R_{av}^{2}\rangle_{\Lambda}$ is modified at lower order in $1/\mathcal{L}$, which suggests instability in the fluctuations that depend on the cosmological constant. We speculate that this instability can provide a potential reason for enforcing a vanishing cosmological constant: non-smooth behavior in the magnitude of curvature fluctuations seems unphysical.
@@ VI Conclusion
In this letter, we have used a baby quantum gravity model based on quantum geometry to compute spacetime curvature fluctations at the ultra-microscopic level. Our results are generally consistent with heuristic arguments regarding the magnitude of spacetime fluctuations and, as such, suggest that a recently proposed resolution [4, 5] of the cosmological constant problem has at least some theoretical merit. Furthermore, the implication of a novel instability in the vacuum in the presence of a perturbatively small cosmological constant leads to additional speculation that, perhaps, the intrinsic cosmological constant must vanish, consistent with recent observations from the DESI collaboration [6].
This work is merely a first step in examining the kinds of physics that can be extracted from baby quantum gravity models based on quantum geometry. The case of Lorentzian $(\mathbb{Z}_{2})^{4}$ should be looked at but is currently too hard to compute, while other currently computable quantum gravity models using quantum geometry [14, 15, 10] are Euclidean rather than Lorentzian. It would also be important to introduce particles to quantify the relative difference in energies associated with gravitational fluctuations and particle fluctuations.

### v2

@@ III Renormalization
The quantities in previous sections were in physical units and for clarity we now label them explicitly as such with the superfix phys. We then work in Planckian units in which we redefine
$\mathcal{L}$ $={\mathcal{L}^{phys}\over G},\quad\varepsilon={\varepsilon^{phys}\over G},\quad\langle k_{0}\rangle_{\mathcal{L}}={\langle k_{0}^{phys}\rangle\over G},$
$\langle R_{av}\rangle_{\mathcal{L}}$ $=G\langle R_{av}^{phys}\rangle,\quad\langle R_{av}^{2}\rangle_{\mathcal{L}}=G^{2}\langle(R_{av}^{2})^{phys}\rangle,$ (18)
where we also write the regulator explicitly in the expectation values. In terms of the rescaled $\mathcal{L},\varepsilon$, we now have
$\langle k_{0}\rangle_{\mathcal{L}}\approx$ $\frac{3\mathcal{L}}{5}+\frac{3i}{160},$
$\langle R_{av}\rangle_{\mathcal{L}}\approx$ $\frac{9i}{8\mathcal{L}^{2}}-\frac{9i(144\sqrt{\pi}+35)}{2240\sqrt{\pi}\mathcal{L}^{5/2}},$
$\langle R_{av}^{2}\rangle_{\mathcal{L}}\approx$ $\frac{9}{64\sqrt{\pi}}\sqrt{\frac{1}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81}{1024\sqrt{\pi}}\sqrt{\frac{1}{\varepsilon^{2}\mathcal{L}^{7}}}\,.$ (19)
Our approach to the renormalization is to solve for the IR regulator $\mathcal{L}$ in terms of a fixed $\langle k_{0}^{phys}\rangle$. Then
$\langle R^{phys}_{av}\rangle=\langle R_{av}\rangle_{\mathcal{L}}\frac{\langle k_{0}\rangle_{\mathcal{L}}}{\langle k_{0}^{phys}\rangle}=\frac{81iG}{200\langle k_{0}^{phys}\rangle^{2}}+\mathcal{O}(G^{3/2})\,,$ (20)
where we replaced $\mathcal{L}$ in terms of $\langle k_{0}^{phys}\rangle$. As expected, any non-vanishing behavior in the averaged Ricci scalar is purely quantum in nature. Similarly replacing $\mathcal{L}$ in $\langle{R_{av}^{2}}\rangle_{\mathcal{L}}$ and converting back to physical quantitites, we find that the renormalized variance of the Ricci scalar takes the form
$\langle(R^{2}_{av})^{phys}\rangle=\frac{81\sqrt{15}}{8000\sqrt{\pi}}\sqrt{\frac{G^{3}}{(\varepsilon^{phys})^{2}\langle k_{0}^{phys}\rangle^{5}}}+\mathcal{O}(G^{5/2})\,.$ (21)
We see that this is a genuine quantum effect and that the behavior (with an adjustment to the coefficient) remains stable under renormalization, namely, a UV divergence still arises as in Section II in terms of renormalised quantities.
@@ IV Physical Interpretation
The construction in Section II provides a more detailed account of the behavior leading to large fluctuations than in [5, 6, 7]. There, a fixed-size region of spacetime is specified and fluctuations are viewed as resonances in that volume, with the magnitude of the fluctuations depending on the scale of the spacetime volume. However, while this heuristic description is a useful starting point, the notion that one begins by considering a fixed volume spacetime is contrary to the spirit [24] of a background-independent theory of quantum gravity. The model provided here using quantum geometry gives a properly background-independent approach that leads to curvature fluctuations, as we have not specified any fixed metric. Instead, we have merely provided a graph structure, which is akin to providing a manifold or differential structure rather than a metric.
Physically, we view this model as approximating the behavior of a “small” subset of the spacetime manifold prior to specification of the metric. By then imposing a quantum geometry on the manifold and minimum and maximum cutoffs on the sizes of the dimensions of that subset, we find that the approximate size of the subset is determined by the expectation value $\langle k_{0}\rangle$, inducing a metric. Unlike the heuristic in [5, 6, 7], the magnitude of curvature fluctuations depends also on the minimum size of the region of spacetime. One way to see this is that derivatives of metric functions depend both on the discrete difference in the metric functions and inversely on the metric functions themselves. If the metric functions are allowed to approach zero, fluctuations can diverge.
Nonetheless, the behavior of this model still reproduces the hierarchy of energy densities required by [4] so long as appropriately physical regulators are introduced. As we intend to model only a “unit” of quantum spacetime in a way that is insensitive to matter field fluctuations, the IR cutoff should be shorter than the scale where Standard Model fields can see the fluctuations. Otherwise, we would already have seen the effects of the discretisation in the Standard Model.
@@
These estimations support the critical assertion in [4] that $\rho_{QG}\gg\rho_{SM}$, implying that for a given region of spacetime the size of the Lorentzian square examined here, the scale factor is oscillatory rather than exponential. Furthermore, taking the IR length scale at $2\times 10^{-21}\text{ m}$, we find from $\langle k_{0}\rangle$ that the average size of the spacetime domain is on the order of $10^{-21}\text{ m}$, meaning that the wavelength of the fluctuations are of the same order. Thus, averaging over even a cubic centimeter of space would mean summing the fluctuations of $10^{56}$ microscopic domains of spacetime, clearly sufficiently many to average out fluctuations [2]. Similarly, the time scale of these fluctuations is on the order of $10^{-30}\text{ s}$, whereas the smallest time yet measured [27] is on the order of $10^{-19}\text{ s}$, meaning that we may average over $10^{10}$ spacetime domains temporally—again, clearly sufficient to average out the fluctuations.
To see an explicit example of the kind of averaging considered here, see [4, Section 5e]. Essentially, one may consider the volume of a given macroscopic region of spacetime by considering the volume of each Planckian subregion (whose geometry is rapidly fluctuating); by adding up the contributions of each subregion and noting that amplitudes of each oscillation do not depend on time, each spatial slice of the macroscopic spacetime volume is shown to be constant. The implication is that, absent any other effects aside from the quantum geometric effects and the Standard Model vacuum fluctuations, the scale factor would be measured to be constant over any macroscopic scale.
While the above computation should be treated as qualitative, it does show that the argument used in [2, 3, 4]—that the gravitational fluctuations dominate the Standard Model energy density when we consider the quantum gravity regime—is robust across many quantum gravitational models, even those as simple as a $1+1$d discrete quantum system as discussed above.
@@ V A Perturbative Cosmological Constant
In [4], it was also assumed that the intrinsic macroscopic cosmological constant of the spacetime vanishes, and that any cosmological constant must arise from quantum behavior. Interestingly, the model constructed in Section II also provides a hint that this constraint is reasonable. To see this, we consider the behavior of curvature correlators in the presence of a small cosmological constant (such that we may keep only terms linear in $\Lambda$). To that end, we perturb the action and the corresponding correlators merely by writing
Z_Λ= 4 ∫_-1^1 ∫_-1^1 ∫_ε^L ∫_ε^L dk  dℓ  dk_0   dℓ_0   k_0 ℓ_0 exp[iG (S_g - 4 Λk_0 ℓ_0) ]
@@
$Z_{\Lambda}\approx\left(1-\tfrac{4i\Lambda}{G}|\langle k_{0}\rangle_{0}|^{2}\right)Z_{0}\,.$ (24)
@@
$=\left(1+\tfrac{4i\Lambda}{G}|\langle k_{0}\rangle_{0}|^{2}\right)\langle\mathcal{O}{\rangle}_{0}-\tfrac{4i\Lambda}{G}\langle k_{0}\ell_{0}\mathcal{O}\rangle_{0}\,,$ (25)
@@
$\langle k_{0}\ell_{0}R_{av}\rangle$ $\approx\frac{iG}{8}+\frac{27i}{2240}\sqrt{\frac{G^{5}}{\mathcal{L}^{3}}},$
$\langle k_{0}\ell_{0}R_{av}^{2}\rangle$ $\approx(0.0079+0.0264\ln G/\epsilon)(G/\mathcal{L})^{3/2}$ (26)
@@
$\langle R_{av}\rangle_{\Lambda}\approx-\frac{28}{25}\Lambda+0.95\Lambda\sqrt{\frac{G}{\mathcal{L}}}\,,$ (27)
@@
$\langle R_{av}^{2}\rangle_{\Lambda}\approx\frac{9}{64\sqrt{\pi}}\sqrt{\frac{G^{3}}{\varepsilon^{2}\mathcal{L}^{5}}}+\frac{81i\Lambda}{400\sqrt{\pi}}\sqrt{\frac{G}{\varepsilon^{2}\mathcal{L}}}.$ (28)
Note that, as expected, the introduction of a cosmological constant induces a non-vanishing Ricci scalar in the large $\mathcal{L}$ limit. (Caution is advised when comparing this expectation value to the Ricci scalar in a classical vacuum universe: the dynamics here are qualitatively different, despite arising from “quantizing” a classical action.) However, $\langle R_{av}^{2}\rangle_{\Lambda}$ is modified at lower order in $1/\mathcal{L}$, which suggests instability in the fluctuations that depend on the cosmological constant. We speculate that this instability can provide a potential reason for enforcing a vanishing cosmological constant: non-smooth behavior in the magnitude of curvature fluctuations seems unphysical.
@@ VI Conclusion
In this letter, we have used an explicit quantum gravity model based on quantum geometry to compute spacetime curvature fluctations at the ultra-microscopic level. Our results are generally consistent with heuristic arguments regarding the magnitude of spacetime fluctuations and, as such, suggest that a recently proposed resolution [2, 3, 4] of the cosmological constant problem has at least some theoretical merit. Furthermore, the implication of a novel instability in the vacuum in the presence of a perturbatively small cosmological constant leads to additional speculation that, perhaps, the intrinsic cosmological constant must vanish, consistent with recent observations from the DESI collaboration [12]. We also found that $\langle R_{av}\rangle$ is imaginary, as expected for a Lorentzian model, but disappears for large $\mathcal{L}$.
This work is merely a first step in examining the kinds of physics that can be extracted from baby quantum gravity models based on quantum geometry. The case of Lorentzian $(\mathbb{Z}_{2})^{4}$ should be looked at but is currently too hard to compute, while other currently computable quantum gravity models using quantum geometry [17, 18, 19] are Euclidean rather than Lorentzian. While our application has been explicitly tied to small scales, one could also use graphs with increasingly many points to model spacetime at large, while scaling all metric ‘square-lengths’ so that the vertices become in some sense ‘closer’ together. Depending on how the scaling does, the noncommutativity of the differential forms and functions can still be present in the limit, with some of the freedom appearing as a deformation parameter (this parameter would not, however, be expected to be $\hbar$ but the Planck scale if the original discreteness was a Planck scale effect). The fact that quantum Riemannian geometry can be specialised to both the graph case and the classical (as well as noncommutative) continuum case means that these limits in principle make sense, although some quantities could still diverge in the limit. How to formulate a Lorentzian graph geometry in general, however, remains unclear, but it may be possible to borrow some ideas here from causal set theory. Doing quantum gravity on such large graphs, although currently too hard to compute, could then be attempted in lieu of continuum quantum gravity. In principle, one could also sum over all bidirected graphs with a fixed number of vertices, which would be akin to considering all possible differential structures on a given topological space, which is a step beyond quantum gravity as normally conceived. It would also be important to introduce particles, and in our specific application above to quantify the relative difference in energies associated with gravitational fluctuations and particle fluctuations.